integration root cos2A/sinA=
 

Grade:12th pass

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{cos2A}{sinA}dA
I = \int \frac{1-2sin^{2}A}{sinA}dA
I = \int cosecAdA - 2\int sinAdA
I = \int cosecAdA - 2(-cosA)
I = \int \frac{1+tan^{2}\frac{A}{2}}{2tan\frac{A}{2}}dA + 2cosA
I = \int \frac{sec^{2}\frac{A}{2}}{2tan\frac{A}{2}}dA + 2cosA
I = log(tan\frac{A}{2})+ 2cosA+c